OpenStudy (anonymous):

find the equation of the line containing the points (-5,8) and (6,-4)

OpenStudy (anonymous):

Change in y over change in x

OpenStudy (anonymous):

That will give you the slope m. Y=ms+b So, to get b (the y intercept) set y=0 and solve for x, that will be b.

OpenStudy (anonymous):

That's where I get confused because there are fractions involved

OpenStudy (anonymous):

Oops! Y=mx+b

OpenStudy (anonymous):

So far I have y=-12/11x - 60/11 + 8 That's where I get stuck.